OpenStudy (anonymous):

if 40% of a given number is 8 then what is 15% of the given number

OpenStudy (phi):

you could use ratios, or you could solve for 100% of the number, then find 15% ratios: \[ \frac{8}{.40}= \frac{x}{.15} \]

OpenStudy (luigi0210):

\[\frac{ 4 }{ 10 }=\frac{ 8 }{ x }\]

OpenStudy (anonymous):

to find that number, (let's represent it by x) \[\frac{ 40 }{ 100 } \times x = 8\] \[x=20\] Now that you know the number, find 15% of that.. \[\frac{ 15 }{ 100 } \times 20 =??\]

OpenStudy (anonymous):

3.0 ?

OpenStudy (anonymous):

So then you multiply .15 by 20 and get 3.0?

OpenStudy (anonymous):

yes...you are right!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

it is 3...

OpenStudy (anonymous):

yw....(: